OpenStudy (anonymous):

Find the common solution. 3x + 2y = 4 2x - y = 5 ( ? , ? )

OpenStudy (anonymous):

x=2,y=-1

jhonyy9 (jhonyy9):

3x+2y=4 --- 3x=4-2(2x-5)=4-4x+10=14-4x --- 7x=14 --- x=14/7 =2 ---x=2 2x-y=5 --- y=2x-5 --- y=2(2) -5 =4-5 =-1 (2,-1)